# Advanced Detection of Source Code Clones via an Ensemble of Unsupervised Similarity Measures

### v1

@@ 4.6 Results
Table 2 shows us the results obtained for the first benchmark dataset, i.e., the small-scale IR-Plag. It can be seen that some unsupervised similarity measures are pretty good on their own if one looks at the f-measure. However, it is striking that ensembles obtain the best results, so we far exceed our weak baseline. Please note that we do not study time consumption issues here as they have been studied previously [28].
@@
Table 3 establishes a comparative for the state-of-the-art published to this dataset. On the one hand, it can be seen that the ensembles can perform better than CodeBERT. The reason is that they perform better without large data for training. This is even though several combinations of different partitions have been evaluated: 60-20-20, 70-15-15, and 80-10-10 on training, validation and test respectively. We are reporting here the best average obtained. On the other hand, GraphCodeBERT matches our best results, although with the disadvantages already discussed throughout this study, so it might make sense to use our solution when working with small datasets.
@@
Table 4 shows us the results of the different unsupervised similarity measures over the BigCloneBench dataset. These results estimate over 25,000 samples (6% of the dataset) because we do not have enough computational power to analyze the complete dataset. We also discard similarity calculations in scenarios that require more than 20 seconds of computation, otherwise the process may require many days of computation. Furthermore, similarity by output analysis cannot be performed in this context due to the high computational requirements of packaging, compiling, executing and comparing several million code samples. Anyway, if desired and if sufficient computational power is available, our solution can be used on the total set.
@@
Table 5 establishes a comparative for the state-of-the-art published to this dataset. For the fine-tuning of CodeBERT and GraphCodeBERT, we rely on the configuration outlined by the authors of the original papers. At the same time, the numbers relative to our strategies are estimated. The reason is that the bagging and boosting strategies have been tested using the seven most promising unsupervised similarity measures from the 415,416 samples of the test partition. These promising measures have been obtained from a preliminary study, please refer to the source code for more information. The reason is, again, not having enough computational resources to operate over the whole search space. Considering the twenty-one approaches should yield better results, albeit only slightly.
@@
Deckard [17] 0.93 0.02 0.03
RtvNN [40] 0.95 0.01 0.01
@@
CDLH [39] 0.92 0.74 0.82
ASTNN [45] 0.92 0.94 0.93
CodeBERT [10] 0.95 0.93 0.94
FA-AST-GMN [38] 0.96 0.94 0.95
TBBCD [43] 0.94 0.96 0.95
GraphCodeBERT[12] 0.95 0.95 0.95
@@ 5 Discussion
Our experiments demonstrate the effectiveness of ensemble learning for assessing code similarity. Although the results are questionable when the training data are abundant, equivalent performance is observed when working with small datasets (467 samples in the case we have studied).
@@ 5.1 Lessons Learned
    Real-world code can be unstructured. Our ensembles of unsupervised methods can manage this noise and variation, particularly relevant in scenarios with little labeled data.
@@
    A human operator cannot easily interpret the Transformers-like architectures on which CodeBERT is based. Moreover, they have a huge carbon footprint when trained [30]. Our approach may be reasonable in situations where these two aspects are critical.
@@ 5.2 Limitations
Second, scalability presents a challenge; as codebases expand in size and complexity, the computational burden of implementing multiple similarity measures increases considerably. This escalation demands integrating more efficient optimization techniques to manage and mitigate the increased processing requirements.
Lastly, while our ensemble method offers theoretical interpretability, its practical use demands a more pragmatic approach. Further work in this direction is crucial to demonstrate how the model identifies similar code fragments, thereby enhancing transparency and usability in practice.
@@ Acknowledgments
The research reported in this paper has been funded by the Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation, and Technology (BMK), the Federal Ministry for Digital and Economic Affairs (BMDW), and the State of Upper Austria in the frame of SCCH, a center in the COMET - Competence Centers for Excellent Technologies Programme managed by Austrian Research Promotion Agency FFG.

### v2

@@ 4.6 Results
Table 1 shows us the results obtained for the first benchmark dataset, i.e., the small-scale IR-Plag. It can be seen that some unsupervised similarity measures are pretty good on their own if one looks at the f-measure. However, it is striking that ensembles obtain the best results, so we far exceed our weak baseline. Please note that we do not study time consumption issues here as they have been studied previously key-martinez-swqd .
@@
Table 2 establishes a comparative for the state-of-the-art published to this dataset. On the one hand, it can be seen that the ensembles can perform better than CodeBERT. The reason is that they perform better without large data for training. This is even though several combinations of different partitions have been evaluated: 60-20-20, 70-15-15, and 80-10-10 on training, validation, and test, respectively. We are reporting here the best average obtained. On the other hand, GraphCodeBERT matches our best results. However, with the disadvantages already discussed throughout this study, it might make sense to use our solution when working with small datasets.
@@
Table 3 shows us the results of the different unsupervised similarity measures over the BigCloneBench dataset. These results estimate over 25,000 samples (6% of the dataset) because we do not have enough computational power to analyze the complete dataset. We also discard similarity calculations in scenarios requiring more than 20 seconds of computation; otherwise, the process may require many days. Furthermore, similarity by output analysis cannot be performed in this context due to the high computational requirements of packaging, compiling, executing, and comparing several million code samples. If desired, our solution can be used on the total set, and sufficient computational power is available.
@@
Table 4 establishes a comparative for the state-of-the-art published to this dataset. For the fine-tuning of CodeBERT and GraphCodeBERT, we rely on the configuration outlined by the authors of the original papers. At the same time, the numbers relative to our strategies are estimated. The reason is that the bagging and boosting strategies have been tested using the seven most promising unsupervised similarity measures from the 415,416 samples of the test partition. These promising measures have been obtained from a preliminary study; please refer to the source code for more information. The reason is, again, not having enough computational resources to operate over the whole search space.
@@
Deckard jiang2007deckard 0.93 0.02 0.03
RtvNN white2016deep 0.95 0.01 0.01
@@
CDLH wei2017supervised 0.92 0.74 0.82
ASTNN zhang2019novel 0.92 0.94 0.93
CodeBERT key-codebert 0.95 0.93 0.94
FA-AST-GMN wang2020detecting 0.96 0.94 0.95
TBBCD yu2019neural 0.94 0.96 0.95
GraphCodeBERTguo2020graphcodebert 0.95 0.95 0.95
@@ 5 Discussion
Our experiments demonstrate the effectiveness of ensemble learning for assessing code similarity. Although the results are questionable when the training data are abundant, equivalent performance is observed when working with small datasets.
@@ 5.1 Lessons Learned
    Real-world code can be unstructured but our ensembles of unsupervised methods can manage this noise and variation, which is particularly relevant in scenarios with little labeled data.
@@
    A human operator cannot easily interpret the Transformers-like architectures on which CodeBERT is based. Moreover, they have a huge carbon footprint when trained key-martinez-jfis . Our approach may be reasonable in situations where these two aspects are critical.
@@ 5.2 Limitations
Second, scalability presents a challenge; as codebases expand in size and complexity, the computational burden of implementing multiple similarity measures increases considerably. This escalation demands the integration of more efficient optimization techniques to manage and mitigate the increased processing requirements.
Lastly, while our ensemble method offers theoretical interpretability, its practical use demands a more pragmatic approach. Further work in this direction is crucial to demonstrate how the model identifies similar code fragments, improving transparency and usability in practice.
@@ Acknowledgments
This research has been funded by the Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation, and Technology (BMK), the Federal Ministry for Digital and Economic Affairs (BMDW), and the State of Upper Austria in the frame of SCCH, a center in the COMET - Competence Centers for Excellent Technologies Programme managed by Austrian Research Promotion Agency FFG.
Data availability and access
The data and the source code to reproduce this research are available at: https://github.com/jorge-martinez-gil/ensemble-codesim. The manuscript is also available as a preprint at https://arxiv.org/abs/2405.02095.